OpenStudy (anonymous):

You are ordering a pizza and soda for the big game. A deep dish pizza costs $10.25, and a two-liter of soda is $2.00. For each touchdown that the Eagles score on Sunday, Vinny's Pizza lowers the cost by $1.50. How many touchdowns will the Eagles have to score for your bill to come to $4.75? Explain how to solve for the missing value using an equation.

OpenStudy (bahrom7893):

all right let me read the question first.. haha

OpenStudy (bahrom7893):

wait lowers the cost of what? pizza or soda? or both?

OpenStudy (anonymous):

just pizza

OpenStudy (bahrom7893):

p = 10.25 s = 2.00 total bill = 10.25 + 2.00 = 12.25

OpenStudy (anonymous):

5 touch downs... to lower the cost of the bill to 4.75

OpenStudy (bahrom7893):

12.25 - 1.5x = 4.75

OpenStudy (bahrom7893):

where x is the number of touchdowns

OpenStudy (anonymous):

5*1.50= 7.50 12.25-7.50 = 4.50

OpenStudy (bahrom7893):

-1.5x = 4.75-12.25

OpenStudy (bahrom7893):

-1.5x = -7.5 x = 7.5/1.5 = 5, i confirm mad's answer

OpenStudy (anonymous):

^.^ yay I love feeling verified

OpenStudy (anonymous):

lololol thanks guys!

OpenStudy (anonymous):

np ^.^

OpenStudy (anonymous):

thanks